OpenStudy (anonymous):

for what value of t is (t+1, 2,t) orthogonal to (1,1,1)? can anyone help please

OpenStudy (amistre64):

when t.v = 0

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

do you know how to do a dot product?

OpenStudy (anonymous):

yes i just dont understand the question some english is second language

OpenStudy (amistre64):

then you should be good to go ;)

OpenStudy (anonymous):

so i got vatue of t=3/1 that mean not orthogonat am i right

OpenStudy (amistre64):

(t+1, 2,t) dot (1,1,1) = (t+1)+(2)+(t) t+1+2+t = 0 , and solve for t

OpenStudy (anonymous):

yes i got t=3/2

OpenStudy (amistre64):

then when t=3/2, the vectors are orthogonal

OpenStudy (amistre64):

i think its a -3/2 tho

OpenStudy (anonymous):

thant is right

OpenStudy (anonymous):

any idea about this \[give the parametric equestion for the following line.a) the line through (2,9,0) and (-1,5,1)\]

OpenStudy (amistre64):

try not to type text in latex

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

a parametric is defined as the point attached to the scaled vector parts

OpenStudy (amistre64):

given a point (x,y,z), and a vector <a,b,c> the parametric is defined as: x = x + at y = y + bt z = z + ct

OpenStudy (anonymous):

but i dont have vector here do i have to use dot product to find v

OpenStudy (amistre64):

i cant read what the information is, can you retype it?

OpenStudy (anonymous):

i dont have the vatue of the vector do have to use dot product to find vector

OpenStudy (amistre64):

are you given 2 points that need to be on the line?

OpenStudy (amistre64):

given point A and point B the vector between them is simply: A-B